# RL-PLUS: Countering Capability Boundary Collapse of LLMs in Reinforcement Learning with Hybrid-policy Optimization

### v2

@@ Abstract
Reinforcement Learning with Verifiable Reward (RLVR) has significantly advanced the complex reasoning abilities of Large Language Models (LLMs). However, it struggles to break through the inherent capability boundaries of the base LLM, due to its essentially on-policy strategy coupled with LLM’s immense action space and sparse reward. Critically, RLVR can lead to the capability boundary collapse, narrowing the LLM’s problem-solving scope. To address this problem, we propose RL-PLUS, a novel hybrid-policy optimization approach for LLMs that synergizes internal exploitation with external data to achieve stronger reasoning capabilities and surpass the boundaries of base models. RL-PLUS integrates two core components, i.e., Multiple Importance Sampling to address for distributional mismatch from external data, and Exploration-Based Advantage Function to guide the model towards high-value, unexplored reasoning paths. We provide both theoretical analysis and extensive experiments to demonstrate the superiority and generalizability of our approach. Compared with existing RLVR methods, RL-PLUS achieves 1) state-of-the-art performance on six math reasoning benchmarks; 2) superior performance on six out-of-distribution reasoning tasks; 3) consistent and significant gains across diverse model families, with average relative improvements up to 69.2%. Moreover, the analysis of Pass@k curves indicates that RL-PLUS effectively resolves the capability boundary collapse problem.00footnotetext: Work done during Yihong Dong and Xue Jiang’s internship at Tongyi Lab. [^1] Our code is available at https://github.com/YihongDong/RL-PLUS.
@@ 1 Introduction
Despite the empirical successes, some work (Havrilla et al., 2024, Shao et al., 2024, Yue et al., 2025a) points out that current RLVR cannot enable LLMs to acquire novel reasoning abilities, but rather simply utilize reasoning patterns already in the base model. As shown in Figure 1(a), Although the pass@1 performance of RLVR-trained models surpasses that of the base model, its pass@128222The pass@k calculates the proportion of problems the model can potentially solve within a finite (k) number of attempts metric is commonly used to gauge a model’s capability boundary. is substantially lower. This trend suggests that the underlying capability distribution of the base model is broader and that existing RLVR can collapse base model’s capability boundary, thus fundamentally limiting the acquisition of new reasoning pathways.
This limitation stems from an essentially challenge when applying RLVR to LLMs: the potential solution space of LLMs is extremely immense with sparse reward that current RLVR techniques cannot effectively guide the model to explore new and unknown pathways, i.e., outward exploration. The challenge is particularly acute in long reasoning tasks where rewards are contingent upon the successful completion of an entire inferential chain. A single erroneous step can nullify the reward for the entire trajectory, thus failing to provide a positive signal for acquiring new knowledge. Consequently, the model is compelled to focus on inward exploitation, meaning that it refines and optimizes the knowledge and reasoning methods it already possesses, which results in a contraction of the model’s exploratory range and a shrinking of its capabilities. This phenomenon not only prevents the model from acquiring new information or abilities that surpass its base model, but also significantly impedes any sustained enhancement of its overall performance.
The ancient educational principle that “If one learns from others but does not think, one will be bewildered. If, on the other hand, one thinks but does not learn from others, one will be in peril” [^3] A principle from the philosopher and educator Confucius. offers a crucial lens through which to view the limitations of current methodologies for enhancing LLM reasoning. Current RLVR can be viewed as the latter case, which excels at “thinking” through inward exploitation but demonstrates an inadequacy outward exploration due to its inherently on-policy strategy coupled with LLM’s immense action space and sparse reward, i.e., hard to continuous “learning” of new knowledge. Conversely, approaches like Supervised Fine-Tuning (SFT) represent the former case, focusing on imitating solutions but failing to internalize the underlying reasoning principles, leading to brittleness when encountering novel problems.
@@ 2.1 Preliminary Knowledge
To accurately assess a model’s true problem-solving capabilities, we utilize the ‘pass@k’ metric (Chen et al., 2021). It measures the probability of obtaining at least one correct answer within $k$ independent samples for a given problem. Unlike mean accuracy (‘pass@1‘), ‘pass@k‘ provides a more comprehensive view of the model’s reasoning potential and is critical for evaluating whether a method expands the set of solvable problems (Yue et al., 2025a).
@@ 3.3 The Composite RL-PLUS Objective
Table 1: Performance of RL-PLUS against against eight recently proposed RLVR methods and four straightforward baselines.
@@ 4 Experimental Results
In this section, we conduct extensive experiments to demonstrate the effectiveness and generalization of RL-PLUS. Detailed experimental setup can be found in Appendix.
@@
As shown in Table 1, RL-PLUS comprehensively outperforms existing RLVR methods across all evaluated applications, achieving SOTA performance. A comparison with several straightforward baselines clearly demonstrates the benefits of RL-PLUS. SFT can be viewed as a means of learning from external knowledge, while GRPO enables the model to explore solutions on its own through reinforcement learning. The combined “SFT+GRPO” approach yields synergistic gains, illustrating the value of integrating both external knowledge and self-exploration. However, the “GRPO w/ SFT Loss” baseline, which simply adds an SFT loss to the RL training, shows a decline in performance. This suggests that effectively merging these two learning paradigms is a non-trivial challenge. RL-PLUS significantly improves upon “SFT+GRPO” by an average of +5.2 points, showcasing a more potent strategy for this integration. Furthermore, when compared to concurrent methods like LUFFY and ReLIFT, which also incorporate external examples into their training process in some form, RL-PLUS also achieves superior performance, which indicates RL-PLUS offers a more effective way for learning from external knowledge.
@@
In Figure 2, we present the training dynamics of our proposed method and baselines on various benchmarks. As illustrated, RL-PLUS consistently outperforms the alternatives in terms of test accuracy and rewards throughout the training process. Notably, RL-PLUS continues to show a clear upward trend in performance even after the baselines have plateaued. We further analyze the changes in actor entropy during training. We observe that directly incorporating external data during rollouts (the green line in Figure 2) leads to an “entropy explosion”, causing the model’s outputs to become chaotic. In contrast, the entropy of the baseline models collapses to zero over the course of training, indicating a loss of exploratory capability. The entropy of RL-PLUS, however, does not diminish to zero, which suggests that our trained model retains a considerable capacity for exploration. Prior research (Cui et al., 2025b) has established that policy performance is achieved at the cost of policy entropy, and the depletion of entropy marks the upper limit of performance. This implies that RL-PLUS still possesses potential for further improvement. Additionally, the response length can reflect the test-time scaling performance of a method. The steadily increasing response length of RL-PLUS is the indicator of a healthy and robust training state. In contrast, while directly incorporating external data also leads to long response lengths, its low accuracy and high policy entropy suggest that this length stems from unproductive exploration rather than meaningful reasoning.
@@ Appendix C Effect of hyperparameter $\gamma$
Our systematic investigation into the hyperparameter $\gamma$ in RL-PLUS, illustrated in Figure 6, reveals two key findings. First, the model demonstrates considerable robustness, as its performance fluctuates only small across the tested range of $\gamma$. Second, RL-PLUS consistently surpasses the GRPO baseline across all math reasoning benchmarks, irrespective of the specific value of $\gamma$. Further analysis highlights a distinct trend: the model uniformly achieves its peak performance when $\gamma$=0.5. This optimal value holds true not only for the Average test score but also across all individual benchmarks, including AMC, Olympiad, AIME, Minerva, and Math. This suggests that an intermediate value for $\gamma$ strikes an effective balance in the model’s learning process. While the model is not highly sensitive to this parameter, the clear peak establishes $\gamma$=0.5 as a strong default, and there is still potential room for improvement with fine-grained tuning of $\gamma$.
@@ Appendix E Experimental Setup
All experiments are conducted on 8 NVIDIA A100 80G GPUs. For our training, we use the dataset from previous work (Yan et al., 2025), which contains 45,000 prompts from OpenR1-Math-220k with correct reasoning trajectories annotated by Deepseek-R1. In implementing the RL algorithm, we leverage the VeRL framework (Sheng et al., 2024). We set the batch size to 128, the mini-batch size to 64, and the maximum training epoch to 2. For each problem, we generate 8 rollouts, with a maximum response length of 8192 tokens. For our approach, we replace one of the model-generated rollouts with a correct reasoning trajectory from training dataset. It is important to note that we ensure all other RL algorithms maintain the same parameter settings as RL-PLUS to guarantee a fair comparison. For the hyperparameter $\gamma$, we set it to 0.5 in all experiments by default. To validate the applicability of RL-PLUS on various base LLMs, we additionally employ LLaMA-3.1-8B-instruct, Deepseek-Math-7B-instruct, Qwen2.5-Math-1.5B, and Qwen2.5-Math-7B as our base models. By default, we use the Qwen2.5-Math-7B model (Yang et al., 2024) as base model in our experiments.
@@
We compare our approach with two categories of baselines, all trained upon the same base model. The first category comprises eight recently proposed RLVR methods, including: 1) SimpleRL (Zeng et al., 2025) and 2) OpenReasoner-Zero (Hu et al., 2025) are two open-source RL implementations that train starting from the base model using rule-based rewards. SimpleRL employs a token-level, length-rectified GRPO algorithm, while OpenReasoner-Zero utilizes the PPO algorithm. 3) PRIME (Cui et al., 2025a) introduces an implicit process reward based on outcome labels during RL. 4) Oat-Zero (Liu et al., 2025b) modifies GRPO algorithm by removing the standard deviation from the advantage computation and eliminating token-level normalization in the policy loss calculation. 5) DAPO (Yu et al., 2025) optimize the GRPO algorithm by introducing four operations: Clip-Higher, Dynamic Sampling, a Token-Level Policy Gradient Loss, and Overlong Reward Shaping. 6) LUFFY (Yan et al., 2025) leverages off-policy reasoning trajectories to augment GRPO. 7) TAPO (Wu et al., 2025) integrates reasoning templates into GRPO sampling process to enhance the model’s internal reasoning capabilities. 8) ReLIFT (Ma et al., 2025) performs RL and SFT alternately during training. The second category consists of four straightforward baselines: 1) SFT, supervised fine-tuning using external reasoning trajectory data. 2) GRPO (Shao et al., 2024), training with GRPO algorithm on question-answer pairs. 3) SFT+GRPO, a common RL cold-start approach that performs SFT before RL training. 4) GRPO w/ SFT Loss, jointly optimizes the GRPO objective and SFT loss during training.

### v3

@@ Abstract
Reinforcement Learning with Verifiable Reward (RLVR) has significantly advanced the complex reasoning abilities of Large Language Models (LLMs). However, it struggles to break through the inherent capability boundaries of the base LLM, due to its essentially on-policy strategy coupled with LLM’s immense action space and sparse reward. Critically, RLVR can lead to the capability boundary collapse, narrowing the LLM’s problem-solving scope. To address this problem, we propose RL-PLUS, a novel hybrid-policy optimization approach for LLMs that synergizes internal exploitation with external data to achieve stronger reasoning capabilities and surpass the boundaries of base models. RL-PLUS integrates two core components, i.e., Multiple Importance Sampling to address distributional mismatch from external data, and Exploration-Based Advantage Function to guide the model towards high-value, unexplored reasoning paths. We provide both theoretical analysis and extensive experiments to demonstrate the superiority and generalizability of our approach. Compared with existing RLVR methods, RL-PLUS achieves 1) state-of-the-art performance on six math reasoning benchmarks; 2) superior performance on six out-of-distribution reasoning tasks; 3) consistent and significant gains across diverse model families, with average relative improvements up to 69.2%. Moreover, the analysis of Pass@k curves indicates that RL-PLUS effectively resolves the capability boundary collapse problem.00footnotetext: Work done during Yihong Dong and Xue Jiang’s internship at Tongyi Lab. [^1] Our code is available at https://github.com/YihongDong/RL-PLUS.
@@ 1 Introduction
Despite the empirical successes, some work (Havrilla et al., 2024, Shao et al., 2024, Yue et al., 2025a) points out that current RLVR cannot enable LLMs to acquire novel reasoning abilities, but rather simply utilize reasoning patterns already in the base model. As shown in Figure 1(a), although the pass@1 performance of RLVR-trained models surpasses that of the base model, its pass@128222The pass@k calculates the proportion of problems the model can potentially solve within a finite (k) number of attempts metric is commonly used to gauge a model’s capability boundary. is substantially lower. This trend suggests that the underlying capability distribution of the base model is broader and that existing RLVR can collapse the base model’s capability boundary, thus fundamentally limiting the acquisition of new reasoning pathways.
This limitation stems from an essential challenge when applying RLVR to LLMs: the potential solution space of LLMs is extremely immense with sparse reward that current RLVR techniques cannot effectively guide the model to explore new and unknown pathways, i.e., outward exploration. The challenge is particularly acute in long reasoning tasks where rewards are contingent upon the successful completion of an entire inferential chain. A single erroneous step can nullify the reward for the entire trajectory, thus failing to provide a positive signal for acquiring new knowledge. Consequently, the model is compelled to focus on inward exploitation, meaning that it refines and optimizes the knowledge and reasoning methods it already possesses, which results in a contraction of the model’s exploratory range and a shrinking of its capabilities. This phenomenon not only prevents the model from acquiring new information or abilities that surpass its base model, but also significantly impedes any sustained enhancement of its overall performance.
The ancient educational principle that “If one learns from others but does not think, one will be bewildered. If, on the other hand, one thinks but does not learn from others, one will be in peril” [^3] A principle from the philosopher and educator Confucius. offers a crucial lens through which to view the limitations of current methodologies for enhancing LLM reasoning. Current RLVR can be viewed as the latter case, which excels at “thinking” through inward exploitation but demonstrates inadequate outward exploration due to its inherently on-policy strategy coupled with LLM’s immense action space and sparse reward, i.e., hard to continuous “learning” of new knowledge. Conversely, approaches like Supervised Fine-Tuning (SFT) represent the former case, focusing on imitating solutions but failing to internalize the underlying reasoning principles, leading to brittleness when encountering novel problems.
@@ 2.1 Preliminary Knowledge
To accurately assess a model’s true problem-solving capabilities, we utilize the pass@k metric (Chen et al., 2021). It measures the probability of obtaining at least one correct answer within $k$ independent samples for a given problem. Unlike mean accuracy (i.e., pass@1), pass@k provides a more comprehensive view of the model’s reasoning potential and is critical for evaluating whether a method expands the set of solvable problems (Yue et al., 2025a).
@@ 3.3 The Composite RL-PLUS Objective
Table 1: Performance of RL-PLUS against other baselines, where the best-performing result for each benchmark is highlighted in bold and the base model is Qwen2.5-Math-7B for all methods.
@@ 4 Experimental Results
In this section, we conduct extensive experiments to demonstrate the effectiveness and generalization of RL-PLUS. Detailed experimental setups can be found in Appendix.
@@
As shown in Table 1, RL-PLUS comprehensively outperforms existing RLVR methods across all evaluated applications, achieving SOTA performance. A comparison with several straightforward baselines clearly demonstrates the benefits of RL-PLUS. SFT can be viewed as a means of learning from external knowledge, while GRPO enables the model to explore solutions on its own through reinforcement learning. The combined “SFT+GRPO” approach yields synergistic gains, illustrating the value of integrating both external knowledge and self-exploration. However, the “GRPO w/ SFT Loss” baseline, which simply adds an SFT loss to the RL training, shows a decline in performance. This suggests that effectively merging these two learning paradigms is a non-trivial challenge. RL-PLUS significantly improves upon “SFT+GRPO” by an average of +5.2 points, showcasing a more potent strategy for this integration. Furthermore, when compared to concurrent methods like LUFFY and ReLIFT, which also incorporate external examples into their training process in some form, RL-PLUS also achieves superior performance, which indicates that RL-PLUS offers a more effective way for learning from external knowledge.
@@
In Figure 2, we present the training dynamics of our proposed method and baselines on various benchmarks. As illustrated, RL-PLUS consistently outperforms the alternatives in terms of test accuracy and rewards throughout the training process. Notably, RL-PLUS continues to show a clear upward trend in performance even after the baselines have plateaued. We further analyze the changes in actor entropy during training. We observe that directly incorporating external data during rollouts (the green line in Figure 2) leads to an “entropy explosion”, causing the model’s outputs to become chaotic. In contrast, the entropy of the baseline models collapses to nearly zero over the course of training, indicating a loss of exploratory capability. The entropy of RL-PLUS, however, does not diminish to zero, which suggests that our trained model retains a considerable capacity for exploration. Prior research (Cui et al., 2025b) has established that policy performance is achieved at the cost of policy entropy, and the depletion of entropy marks the upper limit of performance. This implies that RL-PLUS still possesses potential for further improvement. Additionally, the response length can reflect the test-time scaling performance of a method. The steadily increasing response length of RL-PLUS is the indicator of a healthy and robust training state. In contrast, while directly incorporating external data also leads to long response lengths, its low accuracy and high policy entropy suggest that this length stems from unproductive exploration rather than meaningful reasoning.
@@ Appendix C Effect of hyperparameter $\gamma$
Our systematic investigation into the hyperparameter $\gamma$ in RL-PLUS, illustrated in Figure 6, reveals two key findings. First, the model demonstrates considerable robustness, as its performance fluctuates only small across the tested range of $\gamma$. Second, RL-PLUS consistently surpasses the GRPO baseline across all math reasoning benchmarks, irrespective of the specific value of $\gamma$. Further analysis highlights a distinct trend: the model uniformly achieves its peak performance when $\gamma$=0.5. This optimal value holds not only for the Average test score but also across all individual benchmarks, including AMC, Olympiad, AIME, Minerva, and Math. This suggests that an intermediate value for $\gamma$ strikes an effective balance in the model’s learning process. While the model is not highly sensitive to this parameter, the clear peak establishes $\gamma$=0.5 as a strong default, and there is still potential room for improvement with fine-grained tuning of $\gamma$.
@@ Appendix E Experimental Setup
All experiments are conducted on 8 NVIDIA A100 80G GPUs. By default, we use the Qwen2.5-Math-7B model (Yang et al., 2024) as the base model in our experiments. For our training, we use the dataset from previous work (Yan et al., 2025), which contains 45,000 prompts from OpenR1-Math-220k with correct reasoning trajectories annotated by Deepseek-R1, and change the rope theta of Qwen2.5-Math-7B from 10000 to 40000 and extend the window size to 16384. In implementing the RL algorithm, we leverage the VeRL framework (Sheng et al., 2024). We set the batch size to 128, the mini-batch size to 64, and the maximum training epoch to 2. For each problem, we use 8 rollout trajectories, with a maximum response length of 8192 tokens. For our approach, one of the model-generated rollouts is replaced with a correct reasoning trajectory from the training dataset. It is important to note that we ensure all other RL algorithms maintain the same parameter settings as RL-PLUS to guarantee a fair comparison. For the hyperparameter $\gamma$, we set it to 0.5 in all experiments by default. To validate the applicability of RL-PLUS on various base LLMs, we additionally extend RL-PLUS to other base models, including LLaMA-3.1-8B-instruct, Deepseek-Math-7B-instruct, and Qwen2.5-Math-1.5B.
@@
We compare our approach with two categories of baselines, all trained upon the same base model. The first category comprises eight recently proposed RLVR methods, including: 1) SimpleRL (Zeng et al., 2025) and 2) OpenReasoner-Zero (Hu et al., 2025) are two open-source RL implementations that train starting from the base model using rule-based rewards. SimpleRL employs a token-level, length-rectified GRPO algorithm, while OpenReasoner-Zero utilizes the PPO algorithm. 3) PRIME (Cui et al., 2025a) introduces an implicit process reward based on outcome labels during RL. 4) Oat-Zero (Liu et al., 2025b) modifies the GRPO algorithm by removing the standard deviation from the advantage computation and eliminating token-level normalization in the policy loss calculation. 5) DAPO (Yu et al., 2025) optimizes GRPO algorithm by introducing four operations: Clip-Higher, Dynamic Sampling, a Token-Level Policy Gradient Loss, and Overlong Reward Shaping. 6) LUFFY (Yan et al., 2025) leverages off-policy reasoning trajectories to augment GRPO. 7) TAPO (Wu et al., 2025) integrates reasoning templates into GRPO sampling process to enhance the model’s internal reasoning capabilities. 8) ReLIFT (Ma et al., 2025) performs RL and SFT alternately during training. The second category consists of four straightforward baselines: 1) SFT, supervised fine-tuning using external reasoning trajectory data. 2) GRPO (Shao et al., 2024), training with GRPO algorithm on question-answer pairs. 3) SFT+GRPO, a common RL cold-start approach that performs SFT before RL training. 4) GRPO w/ SFT Loss, jointly optimizes the GRPO objective and SFT loss during training.